OpenStudy (lilmike234):

State the degree rst+r^2s^2+r^3s^2

OpenStudy (freckles):

The degree of the polynomial is going to be highest degree of the terms of the polynomial. For example rst has degree (1+1+1)=3 (since each of the exponents there are 1) can you find the degree of r^2s^2?

OpenStudy (freckles):

hint: just add the exponents as I did for the example I gave.

OpenStudy (lilmike234):

I just did that and came up with 12

OpenStudy (freckles):

2+2=?

OpenStudy (freckles):

the degree of rst is (1+1+1)=3 the degree of r^2s^2 is (2+2)=? (this shouldn't compute to 12)

OpenStudy (freckles):

can you figure out the degree of r^3s^2 ?

OpenStudy (lilmike234):

5

OpenStudy (freckles):

yes because 3+2=5

OpenStudy (freckles):

the degree of rst is 3 the degree of r^2s^2 is 4 the degree of r^3s^2 is 5 the degree of your polynomial will be the highest of those numbers

OpenStudy (lilmike234):

So my degree would be 5 correct ?

OpenStudy (freckles):

yes since 5 is bigger than 3 and 4